For the year ended December 31, 20x5, Davis Corporation has the following records of its costs: Direct materials used $559961 Direct labour 166618 Variable manufacturing overhead 77116 Fixed manufacturing overhead 137224 Selling and administrative costs (fixed) 110477 Davis produced 124569 units and sold 99727 units at $13 in 20x6. There were no beginning inventories. If Davis uses variable costing, what is the operating income for the year ended December 31, 20x6?

Select one:

a. $432696

b. $405331

c. $385331

d. $245055
Answer #1

Unit product cost

= (559961+166618+77116)/124569

= 6.4518

Total Contribution margin = (13-6.4518)*99727

= 653032

Operating income = 653032-(137224+110477)

= 405,331

Option B